OpenStudy (anonymous):

SOMEONE PLEASE HELP: Use the distributive property to simplify the expression below -3x-9/-3

OpenStudy (anonymous):

Multiplication distributes over addition. Division is the same as multiplying by a fraction and subtraction is the addition of a negative number. i.e. \[\frac{-3x-9}{-3} = -\frac{1}{3}(-3x+(-9))\] Multiply everything in parentheses by the -1/3.